# Pairwise vs Higher-order interactions: Can we identify the interaction type in coupled oscillators from time series?

### v1

@@ Introduction
Figure 5: Inferring the pairwise and three-body couplings in 100 oscillators. A. Inferred coupling network. The true coupling network is shown on the left and the inferred coupling from phase time series of 200 cycles by the proposed method is shown on the right. The pairwise and three-body couplings (i.e. individual interaction) are shown separately to illustrate the interaction. Correctly inferred (true positive) pairwise and three-body couplings are shown in black and yellow respectively, while incorrectly inferred (false positive) pairwise and three-body couplings are shown in red and blue respectively. B. Inferred coupling strengths (top: $\hat{W}^{(2)}_{ij}$ and bottom: $\hat{W}^{(3)}_{ijl}$) are plotted against true strengths (top: $W^{(2)}_{ij}$ and bottom: $W^{(3)}_{ijl}$). The points on the non-zero x-axis and y-axis represent the false negatives and false positives, respectively.
@@ Discussion
In this study, we have addressed the question of whether it is possible to distinguish between the types of interaction (i.e. pairwise, three-body, and mixture interactions, all of which contribute to synchronisation) from time series data. As seen in Figure 1, the dynamics of the systems with the three types of interaction were similar, indicating the need for advanced methods to accurately distinguish between them. In order to identify the interaction type, we have developed a method based on the adaptive LASSO for inferring the couplings (i.e. individual interaction) from phase time series. The proposed method is capable of inferring the interaction type and the probability of pairwise and three-body couplings from time series (Tables 1 and 2). Our method has been shown to outperform two baseline methods (LASSO and OLS) in inferring the topology (Figs. 2 and 3, and Table 3) and strength (Fig. 4 and Table 4) of couplings. Finally, we have shown that the proposed method is effective for inferring the couplings in a larger system of oscillators (Fig. 5).
@@
Our study is based on three main assumptions. Firstly, we assume that phase data $\{\phi_{i}(t)\}$ ($i=1,2,...,N$) is available. It should be noted that in many real-world cases, an oscillatory signal is observed instead of the phase; however, the phase can still be reconstructed from the observed signal. In cases where the observed signal is similar to a sinusoidal wave, the Hilbert transform can be used to reconstruct the phase [16, 52, 53, 54]. In cases where marked events such as neural spikes are observed, linear interpolation is often used to reconstruct the phase [29, 30, 31]. Second, we assume that the functional forms of the coupling functions are known: $f^{(2)}_{i}\left(\vec{\phi};{\bf W}^{(2)}_{i}\right)$ and $f^{(3)}_{i}\left(\vec{\phi};{\bf W}^{(3)}_{i}\right)$ (Eqs. 2 and 3). These coupling functions can be derived by applying the phase reduction technique to weakly coupled Stuart-Landau oscillators [20], thereby supporting the validity of the assumption in the vicinity of the Hopf bifurcation. When the oscillators are strongly synchronised, the model with first-order harmonic couplings is capable of reproducing small fluctuations around the synchronised state [44]. However, if the oscillators are not synchronised, coupling functions with higher harmonics would be suitable to describe a global coupling function [41, 42, 43]. Thirdly, we have focused on the pairwise and three-body interactions. It is straightforward to extend the present method to systems with four- or higher-order interactions. However, the number of parameters is $O(N^{k})$ and the required computational cost is $O(N^{3k})$, which depends exponentially on the order $k$ of the interaction. We inferred 10,000 parameters for pairwise interactions and 500,000 parameters for three-body interactions in the case of 100 oscillators (Fig. 5). From a computational perspective, it would be challenging to infer four-body or higher-order interactions from data for a large population (100 or more) of oscillators. Consequently, it is an important future study to develop more efficient algorithms for inferring higher-order interactions.
@@
In this study, we have demonstrated that it is possible to distinguish between pairwise and three-body interactions and to accurately infer the topology and strength of couplings from time series, even when the oscillators are subject to strong noise. Notably, the interactions obtained here, in contrast to the simple correlation, can characterise the causal relationship between the oscillators. A promising future direction is the application of the developed method to data measured from human cardiorespiratory system, brain, etc. The inferred interaction is expected to serve as a potential biomarker for disease [55], and we will refine the network inference method so that it can be applied to more complex oscillatory systems.
@@ Methods
where, $\mu$ and $\sigma^{2}$ are the mean and variance of the natural logarithm of $x$. The parameters are $\mu=0.01$ and $\sigma^{2}=0.25$ for pairwise and three-body coupling strength in Fig. 4 and Fig. 5, respectively. The number of oscillators was $N=10$ in Figs. 1, 3, and 4 and Table 3, $N=12$ in Fig. 2 and Tables 1 and 2, and $N=100$ in Fig. 5.
@@
We describe the baseline methods, i) naive OLS, ii) OLS with statistical test, and iii) LASSO, which are compared with the proposed method for inferring the couplings (i.e. individual interaction) from time series.
@@
As a more refined method, we identify the existence of couplings by using the Benjamini-Hochberg procedure [56, 57] to address the multiple testing problem. Specifically, we obtain the $p$-values $P^{(2)}_{ij},P^{(3)}_{ijl}$ for the coupling strength $\hat{W}^{(2)}_{ij},\hat{W}^{(3)}_{ijl}$ obtained from the original data by assuming that the inferred strength follows a Gaussian distribution under the null hypothesis. We sort the $p$-values $P^{(2)}_{ij},P^{(3)}_{ijl}$ in ascending order, and denote them by $\tilde{P}^{(2)}_{1},\tilde{P}^{(2)}_{2},\cdots\tilde{P}^{(2)}_{N(N-1)}$ and $\tilde{P}^{(3)}_{1},\tilde{P}^{(3)}_{2},\cdots\tilde{P}^{(3)}_{N(N-1)(N-2)/2}$. Then, we find the largest indices $k^{(2)}$ and $k^{(3)}$ such that $\tilde{P}^{(2)}_{k^{(2)}}\leq\frac{k^{(2)}}{m}\alpha$ and $\tilde{P}^{(3)}_{k^{(3)}}\leq\frac{k^{(3)}}{m}\alpha$, where $\alpha=0.05$ is the significance level. Finally, we reject the null hypothesis corresponding to the $p$-values: $\tilde{P}^{(2)}_{1},\cdots,\tilde{P}^{(2)}_{k^{(2)}}$ and $\tilde{P}^{(3)}_{1},\cdots,\tilde{P}^{(3)}_{k^{(3)}}$. If the null hypothesis is rejected, we detect an interaction ($W^{(2)}_{ij}$ or $W^{(3)}_{ijl}$) and its strength is estimated as $\hat{W}^{(2)}_{ij},\hat{W}^{(3)}_{ijl}$. Otherwise, we infer that there is no interaction among the node pair or triple: $W^{(2)}_{ij}=0$ or $W^{(3)}_{ijl}=0$.
@@
$E_{\rm LAS}=E_{\rm OLS}+\alpha\left\{\sum_{j:j\neq i}\left|W^{(2)}_{ij}\right|+\sum_{j,l:l>j,j\neq i,l\neq i}\left|W^{(3)}_{ijl}\right|\right\},$ (11)
@@
    We test the null-hypothesis of $H_{0}:q^{(2)}=q^{(3)}$ using the $Z$-test. Under the null-hypothesis of $H:q^{(2)}=q^{(3)}$, the difference of the coupling probabilities approximately obeys the Gaussian distribution of mean 00 and variance $\sigma^{2}_{d}=q_{c}(1-q_{c})\left(\frac{1}{n^{(2)}}+\frac{1}{n^{(3)}}\right)$, where $q_{c}:=q^{(2)}=q^{(3)}$ is the overall coupling probability. Thus, we can test the null-hypothesis $H_{0}$ by calculating the standardized statistic:
@@ Acknowledgements
This research was partially supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI (Nos. JP22K11919 and JP22H00516) and JST CREST (No. JPMJCR1913) to H.N., JSPS KAKENHI (Nos. JP21K12056, JP22K18384, and JP23K27487) to H.K., and JSPS KAKENHI (Nos. JP18K11560, JP19H01133, JP21H03559, JP21H04571, and JP22H03695), JST FOREST (No. JPMJFR232O), and AMED (No. JP223fa627001) to R.K.
Supplementary Information
Supplementary Figure 1. Effect of coupling strength on inference accuracy. Performance of coupling inference (MCC) was calculated for the three types of interaction: A. Pairwise ($q^{(2)}=q$, $q^{(3)}=0$), B. Three-body ($q^{(2)}=0$, $q^{(3)}=q$), and C. Mixture of pairwise and three-body interaction ($q^{(2)}=q/2$, $q^{(3)}=q/2$). The parameter of the coupling probability $q$ was set to 0.05, 0.10 and 0.15 on the left, centre and right panel, respectively.

### v2

@@ Results
Figure 5: Application to a human brain network (Subject ID: S038). (a) Inferred coupling network. The proposed method was applied to infer the couplings of the human brain network from time series data. The true network (left) is compared with the inferred network (right). Top panels display pairwise couplings; bottom panels show three-body couplings. Only strong couplings with magnitudes greater than 0.05 are visualized. Correctly inferred (true positive) couplings are shown in black (top: pairwise) and yellow (bottom: three-body). Incorrectly inferred (false positive) couplings are shown in red (top: pairwise) and blue (bottom: three-body). (b) Inferred coupling strengths. Scatter plots show inferred versus true coupling strengths for pairwise (top) and three-body (bottom) interactions. Points on the non-zero x-axis represent false negatives (missed true couplings), while those on the non-zero y-axis represent false positives (spurious couplings).
Application to a real-world network
As demonstrated in previous subsections, the proposed method accurately infers both the interaction type and the couplings from time series using synthetic data sets with a small number of oscillators (Table 1, Figs. 2 and 3: 12 oscillators). To further assess the feasibility of our method, we apply it to reconstruct the human brain network. We use the human brain structural connectivity dataset [52] obtained from healthy subjects. The network nodes correspond to 90 cortical regions of interest (ROIs), as defined by the Automatic Anatomical Labeling (AAL) atlas. Structural connectivity matrices are derived from probabilistic tractography applied to diffusion MRI data. The human brain networks are constructed as follows: the link from the node $i$ to $j$ is connected when the structural connectivity from $i$ to $j$ is at least $0.05$. This network is directed because the structural connectivity matrices were asymmetric. Pairwise coupling strengths $W^{(2)}_{ij}$ are assigned to the links using a truncated lognormal distribution, which mimics the distribution observed in neuronal circuits [12, 23]. All values are ensured to be greater than $0.05$. Three-body couplings are assigned to triplets of nodes where pairwise links exist between each pair (i.e., between nodes $i$ and $j$, $j$ and $l$, and $l$ and $i$). The direction of each three-body coupling is determined randomly with equal probability. Here, only the three-body couplings were added, while the pairwise couplings remained unchanged. Three-body coupling strengths $W^{(3)}_{ijl}$ are also drawn from a truncated lognormal distribution, with all values exceeding $0.05$. Finally, phase time series are generated by simulating the coupled oscillators (1) using the pairwise and three-body coupling strengths derived from the human brain data. The inference is performed using 200 cycles of simulated time series.
Figure 5A shows the inference results for strong couplings of a human brain network (Subject ID: S038). While our method detects some false positives in pairwise interactions even for strong couplings, it does not detect any false positives in three-body interactions. The overall MCC was 0.694 with 90 false positives and 102 false negatives. Furthermore, our method accurately estimates the coupling strengths (Fig. 5B). The mean squared error $\epsilon^{2}$ of the inferred coupling strengths across all couplings was $1.15\times 10^{-6}$. Table 5 compares the inference performance on the human brain networks (four subjects) across three methods: our proposed method, LASSO, and OLS. Our method consistently outperforms the approaches in reconstructing the coupling structure. These results highlight the potential of our method for inferring real-world network structures from time series.
Extensions of the proposed method
We have examined the coupling inference problem for a simple oscillator system: Kuramoto-type coupled oscillators with both pairwise and three-body interactions. While the Kuramoto-type model can qualitatively explain various synchronous phenomena [18, 1], the coupling function of weakly coupled oscillators can be described as a sum of basis functions [28, 38]. In this subsection, we further extend the proposed method to more general oscillator systems.
We begin with a system of oscillators exhibiting phase-dependent amplitude [8], capable of describing a wide range of real-world oscillatory systems. Assuming symmetric three-body couplings, a weakly coupled oscillator system can be approximately described by the following equation (see Supplementary Note 2 for details):
$\frac{d\phi_{i}}{dt}=\omega_{i}+f^{(2)}_{i}\left(\vec{\phi};{\bf W}^{(2)}_{i}\right)+f^{(3,1)}_{i}\left(\vec{\phi};{\bf W}^{(3,1)}_{i}\right)+f^{(3,2)}_{i}\left(\vec{\phi};{\bf W}^{(3,2)}_{i}\right)+\sigma_{i}\xi_{i}(t),$ (7)
where $\vec{\phi}=(\phi_{1},\cdots,\phi_{N})$ is the vector of oscillator phases and ${\bf W}^{(2)}_{i}=(W^{(2)}_{i1},\cdots,W^{(2)}_{iN})$, ${\bf W}^{(3,1)}_{i}=(W^{(3,1)}_{i12},\cdots,W^{(3,1)}_{i1N},W^{(3,1)}_{i23}$$\cdots,$ $W^{(3,1)}_{iN-1N})$, and ${\bf W}^{(3,2)}_{i}=(W^{(3,2)}_{i12},\cdots,W^{(3,2)}_{i1N},W^{(3,2)}_{i23},\cdots,W^{(3,2)}_{iN-1N})$ are the vectors of pairwise and three-body coupling strength, respectively. Using the Brent method [35], the inference problem for coupled oscillators with phase-dependent amplitude can be reduced to the estimation problem described by Eq. (7) (see Supplementary Note 2). To evaluate performance, we apply the proposed framework to a small oscillator system ($N=12$) with coupling probabilities: $q^{(2)}=6.8\%$, $q^{(3,1)}=2.0\%$, and $q^{(3,2)}=1.1\%$. The resulting MCC of $0.899$ – substantially higher than those obtained by baseline methods (LASSO: 0.807; OLS: 0.702) – demonstrates the effectiveness of our framework in inferring coupling structures in a broader class of oscillatory systems.
Next, we investigate whether the proposed framework can infer higher-order interactions beyond triadic couplings. Specifically, we consider a system that incorporates four-body interactions (see Supplementary Note 3 for details):
$\frac{d\phi_{i}}{dt}=\omega_{i}+f^{(2)}_{i}\left(\vec{\phi};{\bf W}^{(2)}_{i}\right)+f^{(3)}_{i}\left(\vec{\phi};{\bf W}^{(3)}_{i}\right)+f^{(4)}_{i}\left(\vec{\phi};{\bf W}^{(4)}_{i}\right)+\sigma_{i}\xi_{i}(t),$ (8)
where $\vec{\phi}=(\phi_{1},\cdots,\phi_{N})$ is the vector of oscillator phases and ${\bf W}^{(2)}_{i}=(W^{(2)}_{i1},\cdots,W^{(2)}_{iN})$, ${\bf W}^{(3)}_{i}=(W^{(3)}_{i12},\cdots,W^{(3)}_{i1N},W^{(3)}_{i23},\cdots,W^{(3)}_{iN-1N})$, and ${\bf W}^{(4)}_{i}=(W^{(4)}_{i123},\cdots,W^{(4)}_{i12N},W^{(4)}_{i134},\cdots,W^{(4)}_{iN-2N-1N})$ are the vectors of pairwise, three-body, and four-body coupling strength, respectively. As with the three-body case, our method can be extended to handle four-body interactions (see Supplementary Note 3). To assess performance, we apply the method to a system of $N=12$ oscillators with coupling probabilities $q^{(2)}=5.3\%$, $q^{(3)}=1.4\%$, and $q^{(4)}=0.5\%$. The resulting MCC of $0.782$ outperforms the baseline methods (LASSO: $0.577$; OLS: $0.307$). This result demonstrates that our framework is capable of inferring four-body interactions. However, the inference performance for four-body couplings is lower than that for pairwise and three-body interactions, suggesting that higher-order inference is more challenging and may require longer time series.
@@ Discussion
In this study, we have addressed the question of whether it is possible to distinguish between the types of interaction (i.e. pairwise, three-body, and mixture interactions, all of which contribute to synchronisation) from time series data. As seen in Figure 1, the dynamics of the systems with the three types of interaction were similar, indicating the need for advanced methods to accurately distinguish between them. In order to identify the interaction type, we have developed a method based on the adaptive LASSO for inferring the couplings (i.e. individual interaction) from phase time series. The proposed method is capable of inferring the interaction type and the probability of pairwise and three-body couplings from time series (Tables 1 and 2). Our method has been shown to outperform two baseline methods (LASSO and OLS) in inferring the topology (Figs. 2 and 3, and Table 3) and strength (Fig. 4 and Table 4) of couplings. Furthermore, we have demonstrated the practical utility of the proposed method through its application to human brain networks (Fig. 5 and Table 5). Finally, we have demonstrated that the proposed method can be extended to more general oscillatory systems, including those exhibiting limit cycle deformations and systems with four-body interactions.
@@
Our study is based on three main assumptions. Firstly, we assume that phase data $\{\phi_{i}(t)\}$ (i=1,2,,Ni=1,2,...,N) is available. It should be noted that in many real-world cases, an oscillatory signal is observed instead of the phase; however, the phase can still be reconstructed from the observed signal. In cases where the observed signal is similar to a sinusoidal wave, the Hilbert transform can be used to reconstruct the phase [46, 26, 19, 34]. In cases where marked events such as neural spikes are observed, linear interpolation is often used to reconstruct the phase [18, 60, 37]. Second, we assume that the functional forms of the coupling functions are known: $f^{(2)}_{i}\left(\vec{\phi};{\bf W}^{(2)}_{i}\right)$ and $f^{(3)}_{i}\left(\vec{\phi};{\bf W}^{(3)}_{i}\right)$ (Eqs. 2 and 3). These coupling functions can be derived by applying the phase reduction technique to weakly coupled Stuart-Landau oscillators [30], thereby supporting the validity of the assumption in the vicinity of the Hopf bifurcation. When the oscillators are strongly synchronised, the model with first-order harmonic couplings is capable of reproducing small fluctuations around the synchronised state [35]. However, if the oscillators are not synchronised, coupling functions with higher harmonics would be suitable to describe a global coupling function [45, 43, 42]. Thirdly, we have focused on the pairwise and three-body interactions. It is straightforward to extend the present method to systems with four- or higher-order interactions. However, the number of parameters is $O(N^{k})$ and the required computational cost is $O(N^{3k})$, which depends exponentially on the order $k$ of the interaction. In the case of 90 oscillators, we estimated 8,000 parameters for pairwise interactions and 352,000 parameters for three-body interactions (Fig. 5). From a computational perspective, it would be challenging to infer four-body or higher-order interactions from data for a large population (100 or more) of oscillators. Consequently, it is an important future study to develop more efficient algorithms for inferring higher-order interactions. Another direction for future research is the development of an efficient inference algorithm that focuses on global measures of the interaction network, such as the eigenvalues and eigenvectors of its Laplacian matrix.
@@
In this study, we have demonstrated that it is possible to distinguish between pairwise and three-body interactions and to accurately infer the topology and strength of couplings from time series data, even in the presence of strong noise. Notably, the inferred interactions can characterize causal relationships between oscillators. One promising avenue for future research would be to apply the proposed method to physiological data from cardiorespiratory or neuronal systems. The inferred interactions could serve as potential disease biomarkers [21]. While the current study focuses on the weakly synchronized regime, extending the method to accommodate highly synchronous systems [35] remains an important challenge for future work.
@@ Methods
where, $\mu$ and $\sigma^{2}$ are the mean and variance of the natural logarithm of $x$. The parameters are $\mu=0.01$ and $\sigma^{2}=0.25$ for pairwise and three-body coupling strength in Fig. 4 and Fig. 5, respectively. With the exception of Fig. 5, the number of oscillators is $N=12$. In Fig. 5, the number of oscillators is $N=90$.
@@
We describe the baseline methods: (i) naive ordinary least squares (OLS), (ii) OLS with statistical testing, and (iii) LASSO. The performance of these methods in inferring the couplings is compared with that of the proposed method.
@@
[…]
@@
[…]
@@
    We test the null-hypothesis of $H_{0}:q^{(2)}=q^{(3)}$ using the $Z$-test. Under the null-hypothesis of $H:q^{(2)}=q^{(3)}$, the difference of the coupling probabilities approximately obeys the Gaussian distribution of mean 0 and variance $\sigma^{2}_{d}=q_{c}(1-q_{c})\left(\frac{1}{n^{(2)}}+\frac{1}{n^{(3)}}\right)$, where $q_{c}:=q^{(2)}=q^{(3)}$ is the overall coupling probability. Thus, we can test the null-hypothesis $H_{0}$ by calculating the standardized statistic:
@@ Acknowledgements
We thank Riccardo Muolo for the constructive comments on the manuscript. This research was partially supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI (Nos. JP22K11919 and JP22H00516) and JST CREST (No. JPMJCR1913) to H.N., JSPS KAKENHI (Nos. JP21K12056, JP22K18384, and JP23K27487) to H.K., and JSPS KAKENHI (Nos. JP18K11560, JP19H01133, JP21H03559, JP21H04571, and JP22H03695), JST FOREST (No. JPMJFR232O), and AMED (No. JP223fa627001) to R.K.
Author contributions
R.K., S.H. and H.N conceived the project. W.S. and R.K. developed the methods for inferring the network structure from time series. W.S. performed numerical simulations and analyzed data. W.S., S.H. and R.K. created the figures. R.K. and H.K supervised the project. R.K wrote the draft based on input from W.S., S.H. and H.N. All authors edited and approved the manuscript.
Competing interests
The authors declare no competing interests.
Data availability
The datasets and simulation codes for generating the data are available on GitHub: https://github.com/Weiwei-Su112/Adaptive-LASSO-for-higher-order-oscillatory-systems
Code availability
The code for the adaptive LASSO method for inferring the network from phase time series is available on GitHub https://github.com/Weiwei-Su112/Adaptive-LASSO-for-higher-order-oscillatory-systems